OpenStudy (anonymous):

The area of a rectangle can be found by multiplying the length by the width. A rectangular postage stamp has a length of 9/10 inch and a width of 5/6 inch. What is the area of the stamp, in square inches? 3/4 square inches one-seventeenth square inches one-half square inches one and one-fourth square inches

OpenStudy (helder_edwin):

you said it, the area is \[ \large \frac{9}{10}\cdot\frac{5}{6}= \]

OpenStudy (anonymous):

yeah

OpenStudy (helder_edwin):

what do u get?

OpenStudy (anonymous):

i think i know what to do... im just not sure.... so i time it and it equaled 45/60 then i reduced it and 9/12.... but thats not in the optoons

OpenStudy (anonymous):

*thought

OpenStudy (anonymous):

do i have to reduce agaij?

OpenStudy (anonymous):

*agian

OpenStudy (anonymous):

oh never mind i got the answer :) thanks for the help tho

OpenStudy (helder_edwin):

3/4